# Bidirectional Decoding: Improving Action Chunking via Closed-Loop Resampling

### v1

@@ Abstract
Predicting and executing a sequence of actions without intermediate replanning, known as action chunking, is increasingly used in robot learning from human demonstrations. However, its effects on learned policies remain puzzling: some studies highlight its importance for achieving strong performance, while others observe detrimental effects. In this paper, we first dissect the role of action chunking by analyzing the divergence between the learner and the demonstrator. We find that longer action chunks enable a policy to better capture temporal dependencies by taking into account more past states and actions within the chunk. However, this advantage comes at the cost of exacerbating errors in stochastic environments due to fewer observations of recent states. To address this, we propose Bidirectional Decoding (BID), a test-time inference algorithm that bridges action chunking with closed-loop operations. BID samples multiple predictions at each time step and searches for the optimal one based on two criteria: (i) backward coherence, which favors samples aligned with previous decisions, (ii) forward contrast, which favors samples close to outputs of a stronger policy and distant from those of a weaker policy. By coupling decisions within and across action chunks, BID enhances temporal consistency over extended sequences while enabling adaptive replanning in stochastic environments. Experimental results show that BID substantially outperforms conventional closed-loop operations of two state-of-the-art generative policies across seven simulation benchmarks and two real-world tasks.
@@ 1 Introduction
The increasing availability of human demonstrations has spurred renewed interest in behavioral cloning [1, 2]. In particular, recent studies have highlighted the potential of learning from large-scale demonstrations to acquire a variety of complex skills [3, 4, 5, 6, 7, 8]. However, this approach still struggles with two common properties of human demonstrations: (i) strong temporal dependencies across multiple steps, such as idle pauses [4] and latent strategies [9, 10], (ii) large style variability across different demonstrations, including differences in proficiency [11] and preference [12]. Oftentimes, both properties are prevalent yet unlabeled in collected data, posing significant challenges to traditional behavioral cloning, which typically learns a discriminative model to map an input state to a target action.
@@
In response to these challenges, recent works have pursued a generative approach characterized by two key elements: (i) predicting a sequence of actions over multiple time steps and executing all or part of the sequence, known as action chunking [3] or receding horizon [4]; (ii) modeling the distribution of action chunks and sampling from the learned model in an independent [4, 13] or weakly dependent [14, 3] manner during deployment. Some studies find these elements crucial for learning a performant policy in controlled laboratory scenarios [3, 4], while other recent work reports opposite outcomes under practical conditions [6]. The reasons behind these conflicting results remain unclear.
@@
In this paper, we first dissect the influence of action chunking by examining the divergence between learned policies and human demonstrations. We find that when the input contains no or little history observations – short context lengths [15, 16, 17, 3, 18, 19] – increasing the length of action chunks allows a policy to implicitly condition on more past states, improving its ability to capture the temporal dependencies inherent in demonstrations. However, this advantage comes at the cost of fewer recent state observations, which can be crucial for reacting to unexpected changes in stochastic environments, such as those involving action noise or moving targets. This tradeoff raises a crucial question: How can we preserve the strengths of action chunking without suffering from its limitations in reactivity?
@@
To address this, we introduce Bidirectional Decoding (BID), an inference algorithm that integrates action chunking into closed-loop control. Our main idea is to sample multiple predictions at each time step and selectively search for the most desirable action, as illustrated in Fig. 1. Specifically, BID operates on two decoding criteria: (i) backward coherence, which favors samples that are close to the sequence selected at the previous step; (ii) forward contrast, which favors samples that are close to the output of a stronger policy and distant from those of a weaker one. By coupling sequential decision-making both within and across action chunks, BID captures strong temporal dependencies over extended periods while ensuring sufficient reactivity in single-step executions.
@@
The main contributions of this paper are twofold: (a) a thorough analysis of action chunking, and (b) a decoding algorithm to improve it. Our diagnostic simulations in a one-dimensional setting validate our theoretical analysis. Our experiments with two state-of-the-art robotic policies across seven simulations and two real-world tasks show that BID outperforms conventional closed-loop operations of action chunking by more than $26\%$ in relative performance. BID is computationally efficient, model-agnostic, and easy to implement, serving as a plug-and-play component to enhance generative behavior cloning at test time.
@@
Figure 1: Illustration of action chunking in a stochastic environment, where a robot is challenged to catch a box on a moving trolley. (a) Vanilla action chunking [3] executes actions based on previous predictions, resulting in delayed reactions to the latest box location. (b) Receding horizon [4] allows for faster reactions, but leads to a jittery trajectory in the presence of multimodal demonstrations (e.g., both left- and right-handers). (c) Our Bidirectional Decoding explicitly searches for the optimal action from multiple predictions sampled at each time step, achieving long-range consistency while maintaining closed-loop reactivity.
@@ 2 Related Work
Behavioral Cloning. Learning from human demonstrations is becoming increasingly popular in robot learning due to recent advances in robotic teleoperation interfaces [20, 3, 21, 22]. Generative Behavior cloning, which models the distribution of demonstrations, is particularly appealing due to its algorithmic simplicity and empirical efficacy [23, 24, 25, 26, 3, 22, 17]. However, a significant limitation is compounding errors, where deviations from the training distribution accumulate over time [27, 28]. These errors can be mitigated by gathering expert correction data [27, 29, 30, 31, 32] or injecting noise during data collection [33, 34], but such strategies require additional time and effort from human operators. To address this, recent work proposes predicting a sequence of multiple actions into the future, known as action chunking, which reduces the effective control horizon [35, 3, 36, 37]. By handling sequences of actions, action chunking is also better at handling temporal dependencies in the data, such as idle pauses [38, 4] or multiple styles [39, 12, 40, 11]. However, independently drawn action sequence samples may not preserve the necessary temporal dependencies for smooth and consistent execution. Our work provides a thorough analysis of action chunking and proposes a decoding algorithm to improve it.
@@
Decoding Algorithm. Test-time decoding algorithms have been studied in generative sequence modeling for decades, with renewed attention driven by recent advances in large language modeling (LLM). A prominent approach focuses on leveraging internal metrics, e.g., likelihood scores, to improve the quality of generated sequences. Notable examples include beam search [41, 42], truncated sampling [43, 44, 45], minimum Bayes risk decoding [46, 47], and many others [48, 49]. Another line of research explores the synergy of multiple generative models, such as contrastive decoding [50] and speculative decoding [51], which jointly optimize for quality or efficiency. More recently, several studies have highlighted the potential of guiding the decoding or sampling process through the use of an external discriminative model, such as a classifier [52] or reward model [53]. In the context of robot learning, Huang et al. [54] introduced a framework to guide LLM decoding for long-horizon robotic planning. Similarly, Xu et al. [55] proposed the guidance of diffusion models for manipulator design. To the best of our knowledge, our work is the first to explore decoding algorithms for low-level robotic policy. We propose a decoding strategy centered on forward and backward temporal consistency to address the inherent tradeoffs in action chunking.
@@ 3.1 Preliminaries
Consider a dataset of demonstrations $\mathcal{D}=\{\tau_{i}\}_{i=1}^{N}$, where each demonstration $\tau_{i}$ consists of a sequence of state-action pairs $\tau_{i}=\{(s_{1},a_{1}),(s_{2},a_{2}),\cdots,(s_{T},a_{T})\}$ provided by a human expert. These demonstrations often exhibit strong temporal dependencies: an action $a_{t}$ is not only dependent on the current state $s_{t}$, but also influenced by the previous $k$ steps of states and actions $(s_{t-1},a_{t-1},\cdots,s_{t-k},a_{t-k})$ due to unobservable latent variables. Some latent variables can globally influence an entire sequence (e.g., speed, left- vs right-handed), while others may locally influence specific segments within a sequence (e.g., planning strategies). They can also vary significantly between different demonstrations or between different segments of the same sequence. Fig. 3 illustrates the decision process of a human expert, highlighting the inherent temporal dependencies.
To model these temporal dependencies, recent behavior cloning methods have focused on learning the joint distribution of future actions conditioned on past states $\pi(a_{t},a_{t+1},\cdots,a_{t+l}|s_{t-c},\cdots,s_{t})$, or more succinctly $\pi(a_{t:t+l}|s_{t-c:t})$. Here, $c$ denotes the number of past steps included as state inputs, and $l$ represents the number of future steps for action outputs. Training such policies typically involves minimizing the divergence between the model distribution and the data distribution,
@@
Upon training completion, the policy is deployed by sampling a sequence of actions and executing a subset or the entire sequence for $h\in[1,l]$ time steps without re-planning. This approach, commonly referred to as action chunking [3], essentially takes in $c$ states as context and executes $h$ actions. We thus call it a $(c,h)$-policy.
Interestingly, recent works have shown that the choices of context length $c$ and action horizon $h$ play a crucial role in the empirical success of generative behavior cloning [3, 4, 6]. Specifically, extending the context length $c$ does not always improve performance, especially when human demonstrations are limited (refer to Sec. A.2 for more details). Instead, extending the action horizon $h$ has become a common practice in the design of modern policies.
However, a $(c,h)$-policy built with a short context length and a long action horizon stands in stark contrast to human experts, which often consider a longer history while re-planning at each time step, effectively following a $(k,1)$-policy. In the next section, we will analyze why, despite this difference, extending the action horizon can sometimes improve the learned policy in certain scenarios while limiting it in others.
@@
Figure 2: Illustration of the expert decision process, where each action is influenced by multiple past states due to latent variables.
@@ 3.2 Analysis
To understand the influence of action chunking, we focus on the last time step of an action chunk, where the discrepancy between the expert policy and the learned policy is most pronounced. At this time step $t$, the expert, which is a $(k,1)$-policy written as $\pi^{*}\coloneqq\pi^{*}(a_{t}|s_{t-k:t},z_{t-k:t})$, predicts $a_{t}$ by conditioning on $k$ steps of the past states and the corresponding latent variables. In contrast, a learned $(c,h)$-policy, written as $\pi_{(c,h)}\coloneqq\pi_{(c,h)}(a_{t}\mid s_{t-h-c:t-h},a_{t-h:t-1})$ is constrained to observe $c$ steps of the past states and its predicted actions over the past $h-1$ steps.
Considering that recent policies often use a short context length $c$ and a moderate action horizon $h$, we assume the range of temporal dependency modeled by a $(c,h)$-policy is limited:
@@
Under these assumptions, the divergence between a learned policy and an expert policy is attributed to two factors: (i) the importance of unobserved states in predicting the current action, and (ii) the difficulty of inferring them based on the available information. To more clearly see the influence of action horizon on these factors, we next compare the performance of two policies that have the same context lengths but different action horizons, $\pi_{h}\coloneqq\pi_{(c,h)}(a_{t}|s_{t-h-c:t-h},a_{t-h:t-1})$ and $\pi_{h+d}\coloneqq\pi_{(c,h+d)}(a_{t}|s_{t-h-d-c:t-h-d},a_{t-h-d:t-1})$, where $d>0$ is an extended action horizon.
As illustrated in Fig. 3, each policy has access to unique information that is unavailable to the other. $\pi_{h}$ observes some recent states, where $\pi_{h+d}$ is only aware of the executed actions. On the other hand, $\pi_{h+d}$ has access to some earlier states and actions, which precede all information available to $\pi_{h}$. We characterize the importance of observations as follows (formal definitions in Sec. C.1):
@@
The difficulty of inferring each unobserved state hinges on both the relevant observations as well as the environmental stochasticity. We quantify this difficulty as follows (formal definitions in Sec. C.1):
@@
Given that the forward inference is generally easier than the backward inference, the performance difference between $\pi_{h}$ and $\pi_{h+d}$ is bounded by the following (proofs are deferred to Appendix C)
@@
Let $\mathcal{L}$ be a non-linear, convex loss function. Let $\mathcal{S}^{+}\subset\{s_{t-k:t}\}$ be the states both the $(c,h)$ and the $(c,h+d)$ policies observe and let $\mathcal{S}^{-}\coloneqq\{s_{t-k:t}\}\setminus\mathcal{S}^{+}$. Let $C\coloneqq\{a_{t-h-d:t-1}\}\cup\mathcal{S}^{+}$, $G\coloneqq\{a_{t},z_{t-k:t}\}\cup\mathcal{S}^{-}$. Then, we can bound the expected loss of the $(c,h+d)$-policy and the $(c,h)$-policy as:
$\alpha_{f}-\epsilon_{b}(1-P_{b}^{2d})\leq\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]-\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right]\leq-\alpha_{b}+\epsilon_{f}(1-P_{f}^{2d})$ (2)
@@
We next examine two specific environmental settings: highly deterministic and highly stochastic.
In highly deterministic environments, while both policies need to infer the same number of unobserved states, $\pi_{h+d}$ benefits from conditioning on additional actions, which may significantly aid in inferring the corresponding states through its action chunk. If the maximum errors $\epsilon_{f}$ arising from inferring these states are bounded, $\pi_{h+d}$ becomes strictly advantageous:
@@
In a highly deterministic environment, if $a_{t}$ is temporally dependent on at least one state in $\{s_{t-h-c-d:t-h-c-1}\}$ and $\epsilon_{f}$ is finite,
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]<$ $\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right]$ (3)
Conversely, in highly stochastic environments, inferring the unobserved states is challenging, regardless of whether the actions are known. However, the recent states are likely more important than earlier states for predicting the current action $a_{t}$. In this case, $\pi_{h+d}$ becomes strictly more disadvantageous:
@@
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]>$ $\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right]$ (4)
In summary, there is no universally optimal action horizon across all conditions. Determining the appropriate action horizon requires careful consideration of (i) the length of temporal dependencies in the demonstrations and (ii) the level of transition stochasticity present in an environment. When both temporal dependencies and environmental stochasticities are significant, the vanilla action chunking approach leads to an inherent trade-off between these two competing factors.

### v2

@@ Abstract
Predicting and executing a sequence of actions without intermediate replanning, known as action chunking, is increasingly used in robot learning from human demonstrations. Yet, its reported effects on the learned policy are inconsistent: some studies find it crucial for achieving strong results, while others observe decreased performance. In this paper, we first dissect how action chunking impacts the divergence between a learner and a demonstrator. We find that action chunking allows the learner to better capture the temporal dependencies in demonstrations but at the cost of reduced reactivity in stochastic environments. To address this tradeoff, we propose Bidirectional Decoding (BID), a test-time inference algorithm that bridges action chunking with closed-loop operations. BID samples multiple predictions at each time step and searches for the optimal one based on two criteria: (i) backward coherence, which favors samples that align with previous decisions; (ii) forward contrast, which seeks samples of high likelihood for future plans. By coupling decisions within and across action chunks, BID promotes consistency over time while maintaining reactivity to unexpected changes. Experimental results show that BID boosts the performance of two state-of-the-art generative policies across seven simulation benchmarks and two real-world tasks. Code and videos are available at https://bid-robot.github.io.
@@ 1 Introduction
The increasing availability of human demonstrations has spurred renewed interest in behavioral cloning (Atkeson & Schaal, 1997; Argall et al., 2009). In particular, recent studies have highlighted the potential of learning from large-scale demonstrations to acquire a variety of complex skills (Zhao et al., 2023; Chi et al., 2023; Fu et al., 2024b; Lee et al., 2024; Khazatsky et al., 2024). Yet, existing methods still struggle with two common properties of human demonstrations: (i) strong temporal dependencies across multiple steps, such as idle pauses (Chi et al., 2023) and latent strategies (Xie et al., 2021; Ma et al., 2024), (ii) large style variability across different demonstrations, such as differences in proficiency (Belkhale et al., 2024) and preference (Kuefler & Kochenderfer, 2017). Often, both properties are prevalent yet unlabeled in collected data, posing significant challenges to the traditional behavioral cloning that maps an input state to an action.
@@
In response to these challenges, recent works have pursued a generative approach equipped with action chunking: (i) predicting a sequence of actions over multiple time steps and executing all or part of the sequence (Zhao et al., 2023; Chi et al., 2023); (ii) modeling the distribution of action chunks and sampling from the learned model in an independent (Chi et al., 2023; Prasad et al., 2024) or weakly dependent (Janner et al., 2022; Zhao et al., 2023) manner for sequential decisions. Some studies find this approach crucial for learning a performant policy in laboratory scenarios (Zhao et al., 2023; Chi et al., 2023), while other recent work reports opposite outcomes under practical conditions (Lee et al., 2024). The reasons behind these conflicting observations remain unclear.
@@
In this paper, we first dissect the influence of action chunking by examining the divergence between learned policies and human demonstrations. We find that, when a policy is built with limited context length – little or no history is used as input for robustness or efficiency (Mandlekar et al., 2020; Bharadhwaj et al., 2023b; Brohan et al., 2023a; b; Shi et al., 2023; Collaboration et al., 2023) – increasing the length of action chunks allows for implicit conditioning on more past actions, thereby improving its ability to capture the temporal dependencies inherent in demonstrations. However, this advantage comes at the cost of reduced access to recent state observations, which can be crucial for reacting to unexpected changes in stochastic environments, such as action noise or object motions. This tradeoff raises a crucial question: How can we preserve the strengths of action chunking in long-term consistency without suffering from its limitations in short-term reactivity?
@@
To this end, we introduce Bidirectional Decoding (BID), an inference algorithm that bridges action chunking with closed-loop operations. Our main idea is to sample multiple predictions at each time step and search for the most desirable one. Specifically, BID operates on two decoding criteria: (i) backward coherence, which favors samples that are close to the sequence selected at the previous step; (ii) forward contrast, which favors samples that are close to the output of a stronger policy and distant from those of a weaker one. As illustrated in Fig. 1, BID updates the chunk of future actions based on the previous strategy, promoting temporal consistency over extended periods while remaining reactive to unexpected changes.
@@
The main contributions of this paper are twofold: (i) a thorough analysis of action chunking (Sec. 3), and (ii) a decoding algorithm to improve it (Sec. 4). Empirically, we validate our theoretical analysis through a one-dimensional diagnostic simulation and evaluate our decoding method on two state-of-the-art generative policies across seven simulations and two real-world tasks (Sec. 5). Our experiment results show that the proposed BID boosts the performance of recent policies by more than $32\%$ in relative performance. BID is model-agnostic, computationally efficient, and easy to implement, serving as a plug-and-play component to enhance generative behavior cloning at test time.
@@
Figure 1: Illustration of action chunking in a stochastic environment, where a robot is challenged to catch a moving trolley. (a) Vanilla action chunking (Zhao et al., 2023) executes actions based on previous predictions, resulting in delayed reactions to the latest box location. (b) Receding horizon (Chi et al., 2023) allows for faster reactions, but leads to a jittery trajectory in the presence of multimodal demonstrations (e.g., both left- and right-handers). (c) Our Bidirectional Decoding explicitly searches for the optimal action from multiple predictions sampled at each time step, achieving both long-range consistency and short-term reactivity.
@@ 2 Related Work
Behavioral Cloning. Learning from human demonstrations is becoming increasingly popular in robot learning due to recent advances in robotic teleoperation interfaces (Sivakumar et al., 2022; Zhao et al., 2023; Wu et al., 2023; Chi et al., 2024). Generative Behavior cloning, which models the distribution of demonstrations, is particularly appealing due to its algorithmic simplicity and empirical efficacy (Jang et al., 2022; Florence et al., 2022; Brohan et al., 2022; Shafiullah et al., 2022; Zhao et al., 2023; Chi et al., 2024; Brohan et al., 2023a). However, a significant limitation is compounding errors, where deviations from the training distribution accumulate over time (Ross et al., 2011; Ke et al., 2021). These errors can be mitigated by gathering expert correction data (Ross et al., 2011; Kelly et al., 2019; Menda et al., 2019; Hoque et al., 2021a; b) or injecting noise during data collection (Laskey et al., 2017; Brandfonbrener et al., 2023), but such strategies require additional time and effort from human operators. To address this, recent work proposes predicting a sequence of multiple actions into the future, known as action chunking, which reduces the effective control horizon (Lai et al., 2022; Zhao et al., 2023; George & Farimani, 2023; Bharadhwaj et al., 2023a). By handling sequences of actions, action chunking is also better at handling temporal dependencies in the data, such as idle pauses (Swamy et al., 2022; Chi et al., 2023) or multiple styles (Li et al., 2017; Kuefler & Kochenderfer, 2017; Gandhi et al., 2023; Belkhale et al., 2024). However, independently drawn action sequence samples may not preserve the necessary temporal dependencies for smooth and consistent execution. Our work provides a thorough analysis of action chunking and proposes a decoding algorithm to improve it.
@@
Sequential Decoding. Decoding algorithms have been studied in generative sequence modeling for decades, with renewed attention driven by recent advances in large language modeling (LLM). One prominent approach focuses on leveraging internal metrics, e.g., likelihood scores, to improve the quality of generated sequences. Notable examples include beam search (Freitag & Al-Onaizan, 2017; Vijayakumar et al., 2018), truncated sampling (Fan et al., 2018; Hewitt et al., 2022), minimum Bayes risk decoding (Kumar & Byrne, 2004; Müller & Sennrich, 2021), and others (Welleck et al., 2019; Meister et al., 2023; Fu et al., 2024a). Another line of research explores the distinctions between multiple models to jointly optimize for the desired properties such as quality or efficiency (Li et al., 2023; Leviathan et al., 2023). More recently, several studies have highlighted the potential of guiding the decoding or sampling process through the use of an external model, such as a classifier (Dhariwal & Nichol, 2021) or reward model (Khanov et al., 2023). In the context of robot learning, recent works have explored guided decoding for long-horizon robotic planning (Huang et al., 2023) and manipulator geometry designs (Xu et al., 2024). Nevertheless, effective decoding strategies for low-level robotic actions remain lacking. Concurrent to our work, Nakamoto et al. (2024) propose to select action samples by querying a value function learned from reward-annotated demonstrations (Hansen-Estruch et al., 2023). Our work does not rely on a separate value function; instead, we propose a decoding strategy that addresses the consistency-reactivity tradeoff inherent in action chunking through sample comparison.
@@ 3.1 Preliminaries
Consider a dataset of demonstrations $\mathcal{D}=\{\tau_{i}\}_{i=1}^{N}$, where each demonstration $\tau_{i}$ consists of a sequence of state-action pairs $\tau_{i}=\{(s_{1},a_{1}),(s_{2},a_{2}),\cdots,(s_{T},a_{T})\}$ provided by a human expert. At each time step $t$, the demonstrated action $a_{t}$ is influenced not only by the observed state $s_{t}$, but also by latent variables $z_{t}$, such as planning strategies (e.g., subgoals) and personal preferences (e.g., handedness). These latent variables can persist across multiple time steps and vary significantly between different demonstrations. Fig. 3 illustrates the decision process of a human expert, highlighting the inherent temporal dependencies.
To model these temporal dependencies, some recent work (Zhao et al., 2023; Chi et al., 2023; Lee et al., 2024) utilize action chunking, i.e., learning the joint distribution of future actions conditioned on past states $\pi(a_{t},a_{t+1},\cdots,a_{t+l}|s_{t-c},\cdots,s_{t})$, or in short $\pi(a_{t:t+l}|s_{t-c:t})$. Here, $c$ denotes the number of past steps for state inputs, and $l$ denotes the number of future steps for action outputs. Training such a policy typically involves minimizing the divergence of action distributions between the model $\pi$ and the expert $\pi^{*}$,
@@
During deployment, the policy is operated by sampling a sequence of actions and executing a part of or the entire sequence for $h\in[1,l]$ time steps without re-planning. This approach essentially takes in $c$ states as context and executes $h$ actions, which we refer to as a $(c,h)$-policy.
The choices of context length $c$ and action horizon $h$ often play a crucial role in the effectiveness of the learned policy. Recent policies often use a short context length $c$, as extending the context can lead to performance degradation in the presence of limited training data (refer to Sec. A.2 for more details). Conversely, extending the action horizon $h$ has produced mixed results. Some studies report benefits in laboratory settings (Zhao et al., 2023; Chi et al., 2023), while others find it detrimental in real-world scenarios (Lee et al., 2024). The reasons behind these conflicting outcomes are not well understood yet.
Notably, Zhao et al. (2023) hypothesize that action chunking mitigates compounding errors, but it is unclear why this would hold when action chunking cannot correct deviations during open-loop execution. Similarly, Chi et al. (2023) argue that action chunking facilitates consistency and planning, yet the choice of action horizon is rather short and remains an empirical heuristic. The lack of a theoretical understanding of action chunking limits our ability to use it effectively across different methods and tasks. In the following sections, we seek to explicitly identify the strengths and weaknesses of action chunking, aiming to provide practical guidance on its suitable settings as well as to inform the design of better algorithms.
@@
Figure 2: Illustration of the expert decision process, where a latent variable introduces temporal dependencies in actions.
@@ 3.2 Analysis
To understand the influence of action chunking, we focus on the last time step of an action chunk, where the discrepancy between the expert policy and the learned policy is most pronounced. At this time step $t$, a $(k,1)$-expert written as $\pi^{*}\coloneqq\pi^{*}(a_{t}|s_{t-k:t},z_{t-k:t})$ predicts $a_{t}$ by conditioning on $k$ steps of the past states and the corresponding latent variables. In contrast, a $(c,h)$-learner written as $\pi_{(c,h)}\coloneqq\pi_{(c,h)}(a_{t}\mid s_{t-h-c:t-h},a_{t-h:t-1})$ is constrained to observe $c$ steps of the past states and $h-1$ steps of the past actions within the predicted chunk.
Considering that recent policies often use a short context length $c$, we assume the range of temporal dependency modeled by a $(c,h)$-policy is limited:
@@
Under these assumptions, the divergence between a learned policy and an expert policy is attributed to two factors: (i) the importance of unobserved states in predicting the current action, and (ii) the difficulty of inferring unobserved states based on the available information.
To more clearly see the influence of action horizon on these factors, we next compare the performance of two policies that have the same context lengths but different action horizons, $\pi_{h}\coloneqq\pi_{(c,h)}(a_{t}|s_{t-h-c:t-h},a_{t-h:t-1})$ and $\pi_{h+d}\coloneqq\pi_{(c,h+d)}(a_{t}|s_{t-h-d-c:t-h-d},a_{t-h-d:t-1})$, where $d>0$ is the extended action horizon. As illustrated in Fig. 3, each policy has access to unique information that is unavailable to the other. $\pi_{h}$ observes some recent states, where $\pi_{h+d}$ is only aware of the executed actions. On the other hand, $\pi_{h+d}$ has access to some earlier states and actions, which precede all information available to $\pi_{h}$. We characterize the importance of observations as follows (formal definitions in Sec. C.1):
@@
The difficulty of inferring each unobserved state hinges on both the relevant observations as well as the environmental stochasticity. We quantify this difficulty as follows:
@@
Given that the forward inference is generally easier than the backward inference, the performances of $\pi_{h}$ and $\pi_{h+d}$ differ as follows (proofs are deferred to Appendix C):
@@
Let $\mathcal{L}$ be a non-linear convex function measuring the prediction error with respect to demonstrations. Let $\mathcal{S}^{+}\subset\{s_{t-k:t}\}$ be the states both the $(c,h)$ and the $(c,h+d)$ policies observe and let $\mathcal{S}^{-}\coloneqq\{s_{t-k:t}\}\setminus\mathcal{S}^{+}$. Let $C\coloneqq\{a_{t-h-d:t-1}\}\cup\mathcal{S}^{+}$, $G\coloneqq\{a_{t},z_{t-k:t}\}\cup\mathcal{S}^{-}$. Then, we can bound the expected loss of the $(c,h+d)$-policy and the $(c,h)$-policy as:
$\alpha_{f}-\epsilon_{b}(1-P_{b}^{2d})\leq\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]-\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right]\leq-\alpha_{b}+\epsilon_{f}(1-P_{f}^{2d}).$ (2)
@@
We next examine two specific environmental settings: deterministic and stochastic. In highly deterministic environments, while both policies need to infer the same number of unobserved states, $\pi_{h+d}$ benefits from conditioning on additional actions, which may significantly aid in inferring the corresponding states through its action chunk. If the maximum errors $\epsilon_{f}$ arising from inferring these states are bounded, $\pi_{h+d}$ becomes strictly advantageous:
@@
In a highly deterministic environment, if $a_{t}$ is influenced by at least one state in $\{s_{t-h-c-d:t-h-c-1}\}$ and $\epsilon_{f}$ is finite, then
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]<$ $\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right].$ (3)
Conversely, in highly stochastic environments, inferring the unobserved states is challenging, regardless of whether the corresponding actions are known. On the other hand, the recent states are likely more important than earlier states for predicting the current action $a_{t}$. As a result, $\pi_{h+d}$ becomes strictly more disadvantageous:
@@
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]>$ $\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right].$ (4)
In summary, depending on the experimental conditions, action chunking may have varying effects on the learned policy. On the one hand, it benefits the modeling of temporal dependencies in the demonstrations, due to extended access to actions executed at past time steps. On the other hand, it hinders reactions to unexpected states in stochastic environments, due to reduced access to state observations at recent time steps. As a result, there is no universally optimal choice of action horizon across all conditions. When both temporal dependencies and transition stochasticities are significant, tuning the action horizon entails an inherent trade-off between the two opposing factors. Moreover, stochasticitic changes (e.g., wind gusts or object drops) are often unpredictable in advance, raising a critical need to ensure the reactivity of a policy built with action chunking.